OpenStudy (anonymous):

Solve for a. 7a - 2b = 5a + b

OpenStudy (anonymous):

A=1.5b

OpenStudy (anonymous):

I'm sorry not the right answer hold on let me redo the problem

OpenStudy (texaschic101):

7a - 2b = 5a + b -- subtract 5a from both sides 7a - 5a - 2b = b --- add 2b to both sides 7a - 5a = b + 2b -- simplify 2a = 3b -- divide both sides by 2 a = 3/2b

OpenStudy (mrnood):

that should be read as 3b/2 or \[a=\frac{ 3b }{ 2 }\]